OpenStudy (anonymous):

what is 7-9+58x3/2-10+14=?

OpenStudy (anonymous):

7 - 9 + 58 x 1.5 - 10 + 14 is this the question?

OpenStudy (lgbasallote):

three choices: 1) wolframalpha 2) simple/scientific calculator 3) google calculator

OpenStudy (anonymous):

yep. ^ and im too lazy to get a calculator

OpenStudy (anonymous):

I do not believe we are here to give answers, just to assist those in need of guidance and direction get to an answer. However, since you know how to get the answer yourself, I urge that you do the math yourself to receive your answer. @lgbasallote provided a very good compendium of ways to get the answer you need. I would suggest number 1, as it has one of the best calculator algorithms provided free of charge to be accessed by man.

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

\[7-9+58 \times \frac{ 3 }{ 2 }-10+14=?\] you need to use BODMAS rule ; B brackets O of D division M multiplication A addition S subtraction to solve this first b then o then d then m.... you will find the answer ;) should i solve?

OpenStudy (anonymous):

nope got it. thanks:)

Directrix (directrix):

I interpret the problem as: 7 - 9 + 58 x 3/2 - 10 + 14. --------------------------------------------- I don't care for the BODMAS or "Please My Dear Aunt Sally" mnemonic because I do not see either as a heirarchy of operations. For example, Addition and Subtraction are on the same level or importance in the "heirarchy" as are Multiplication and Division. Suggesting to an Asker that he or she crank out 7-9+58x3/2-10+14 via some calculator website is, in my opinion, no different from doling out the answer without explanation on this site. Answers with absolutely no explanations continue to be plastered on this site despite the OS Code of Conduct.Work supporting those answers is sometimes posted with no input from the Asker despite the OS Code of Conduct. Engaging the Asker in mathematical dialogue, even finding the Asker in real time, is tremendously difficult. Patience is tamed passion, and sometimes it is difficult to tame mathematical passion.

OpenStudy (anonymous):

thank youuu. i was not asking for a direct answer(as thats how other viewers of this question perceived i was looking for)however, i was hoping for an explanation on how to acheive this answer.

Directrix (directrix):

7 - 9 + 58 x 3/2 - 10 + 14 is how I interpret the problem. So ... @Kaitlyn_The_Math_wiz

Directrix (directrix):

@Kaitlyn_The_Math_wiz Let's do the problem. Okay?